L & L, CPAs, employs two full-time professional CPAs and five support employees. Budgeted direct salary costs include $160,000 for each CPA. The support employees are considered as indirect costs, and this cost was budgeted for $221,000 although the actual cost was $246,000. Actual salaries were $144,000 for each CPA. Direct and indirect costs are applied on a CPA-labor-hour basis. Total budgeted CPA-labor-hours were 4,000. If a client used 600 labor-hours, what are the budgeted direct-cost rate and the budgeted indirect-cost rate, respectively?
Answer #1

Budgeted direct cost rate = Budgted cost/budgeted labor hours

= 160,000*2/4,000

=$80 per CPA hour

Indirect rate = 221000/4000 =$55.25 per hour

Amount applied is as follows:

Direct =80*600 =$48,000

Indirect = 600*55.25 =$33,150